OpenStudy (anonymous):

(logs)find the solution: 2^(3x)=34

OpenStudy (mertsj):

3xlog2 = log 34

OpenStudy (anonymous):

\[3x=\frac{\ln(34)}{\ln(2)}\] done in one more step

OpenStudy (anonymous):

would it be x=ln(17)/3 ?

OpenStudy (mertsj):

no

OpenStudy (anonymous):

then how do you get x alone?

OpenStudy (mertsj):

Divide the ln of 17 by the ln of 2 and divide the result by 3

OpenStudy (anonymous):

oh okay thank you do much

OpenStudy (mertsj):

yw